half of a 16 in diameter circle fits snugly inside a rectangular box, how many square inches larger is the area of the box than the area of the circle?
Answer by nerdybill(7384)

half of a 16 in diameter circle fits snugly inside a rectangular box, how many square inches larger is the area of the box than the area of the circle?
.
Draw a diagram of the problem -- label everything you know -- it'll help you see the solution.
.
Area of "half circle"
(1/2)(pi)8^2
= (1/2)(pi)8^2
= 32(pi)
= 100.53 sq inches
.
box:
length is 16
width is 8
Area = width*length
Area = 8*16
Area = 128 sq inches
.
It is larger by:
128 - 100.53 = 27.47 sq inches